# Rationale-Enhanced Decoding for Multi-modal Chain-of-Thought

### v1

@@ 1 Introduction
Recent large language model (LLM) advancements exhibit impressive complex reasoning capabilities Achiam_2023_GPT4 ; Google_2023_Gemini ; Bai_2022_Claude . This progress now extends beyond single-modal text-to-text reasoning. The integration of pre-trained vision encoders, like CLIP Radford_ICML21_CLIP , with LLMs has led to the development of large vision-language models (LVLMs) liu_NeurIPS23_llava ; Dai_NeurIPS23_InstructBLIP ; Zhu_ICLR24_MiniGPT ; Chen_NeurIPS24_InternVL ; Bai_2025_Qwen-vl , enabling more complex multi-modal reasoning.
Chain-of-thought (CoT) prompting is a key factor in LLM reasoning abilities Wei_NeurIPS22_CoT ; Kojima_NeurIPS22_lets_think_step_by_step . CoT-prompted models first generate intermediate reasoning steps, termed rationales, then incorporate them into the context to produce the final query response. CoT facilitates models to understand queries deeply and promotes logical and coherent responses Wang_ACL23_towards_understanding_cot ; Saparov_ICLR23_systematic_analysis_cot ; Wang_COLING24_analyzing_cot . This principle has also been applied to multi-modal reasoning in LVLMs Zhang_TMLR_mm-cot ; Lu_NeurIPS22_unifiedqa ; Zheng_NeurIPS23_ddcot ; He_AAAI24_mm_latent_cot ; Mondal_AAAI24_kam-cot ; Mitra_CVPR24_CCoT ; Zheng_LREC24_enhancing_mmcot_soft_negative_sampling ; Zheng_ACMMM24_mm_debate . In CoT for LVLMs, models generate intermediate rationales from the input image and the text query. These rationales, with the original inputs, are then used for grounded image-text-to-text reasoning. Recent research also explores generating structured rationales (e.g., scene graphs) to enhance LVLMs’ capabilities like spatial reasoning Mitra_CVPR24_CCoT ; Zheng_ACMMM24_mm_debate . Thus, similar to text-only models, CoT in pre-trained LVLMs is widely assumed beneficial, offering deeper, multi-modally grounded understanding for more accurate responses.
@@
A straightforward solution to make models force rationale grounding would be to fine-tune LVLMs on datasets containing image-query-rationale-answer tuples, as in Zhang_TMLR_mm-cot ; Zheng_NeurIPS23_ddcot . However, this approach demands expensive annotated datasets and additional training costs. Therefore, we focus on developing a plug-and-play decoding strategy for pre-trained LVLMs. We address the following primary research question: Can we enhance the LVLMs’ grounding capability on the rationales in CoT and improve performance by solely modifying the decoding strategy without additional training?
@@
In this paper, we propose rationale-enhanced decoding (RED), a novel inference-time decoding strategy designed to harmonize information from visual and rationale tokens in LVLMs without additional training. Our core idea is decoupling the next token probability into distinct image-conditional $p(y|x,q)$ and rationale-conditional $p(y|r,q)$ and separately enhancing rationale grounding ($x$: image, $r$: rationale, $q$: query, and $y$: output). Based on this idea, we re-formulate multi-modal CoT prompting as a KL-constrained reward maximization Schulman_ICML15_TRPO ; Schulman_ICML15_PPO . This aims to maximize policy regarding rationale-conditional log-likelihood $\log p(y|r,q)$ as the reward while staying close to image-conditional likelihood $p(y|x,q)$. Solving this problem makes LVLMs explicitly ground on both image and rationale information in the next token prediction. Without additional training, RED provably yields the optimal solution of this maximization by composing the next-token distribution as $p(y|x,q)\times p(y|r,q)^{\lambda}$. Practically, RED is implemented as a simple weighted sum of the log-softmax logits of $p(y|x,q)$ and $p(y|r,q)$, allowing easy, training-free, and plug-and-play integration with existing LVLMs without architecture modification.
@@
We conduct extensive experiments comparing RED against standard CoT prompting and plug-and-play decoding methods for off-the-shelf LVLMs across multiple benchmark datasets and backbone LVLMs. Our results demonstrate that RED consistently and significantly improves reasoning performance. Furthermore, RED’s advantages amplifies with high-quality rationales (e.g., intervening with GPT-4Achiam_2023_GPT4 ). Our findings validate the effectiveness and practicality of RED for enhancing CoT reasoning faithfulness and accuracy in LVLMs, opening avenues for more reliable, interpretable, rationale-grounded multi-modal systems.
@@ 2 Preliminaries
2.1 Next-token Prediction in Large Vision-Language Models
@@
$p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$ $=\operatorname{softmax}(\operatorname{logits}_{\theta}(y_{i}|\bm{y}_{<i},x,q)),$ (2)
@@
Each token $y_{i}$ is generated from $p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$ via a decoding strategy such as greedy decoding, i.e., $y_{i}=\operatorname*{arg\,max}_{w\in\mathcal{V}}p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$. We define $\operatorname{generate(\cdot)}$ as a utility function for an arbitrary decoding strategy:
@@ 2.2 Multi-modal Chain-of-Thought Prompting
Inspired by single modal CoT prompting Wei_NeurIPS22_CoT , multi-modal CoT aims to enhance LVLM reasoning by incorporating the intermediate rationale generation during decoding Zhang_TMLR_mm-cot ; Zheng_NeurIPS23_ddcot ; Mondal_AAAI24_kam-cot ; Mitra_CVPR24_CCoT . In general, multi-modal CoT involves two reasoning steps: (i) rationale generation and (ii) output generation. First, LVLMs generate a rationale $r$ from input image $x$, query prompt $q$ with instruction prompt, e.g., “Given the image, generate the rationale for answering the question”, as follows.
@@
In this line of work, MM-CoT Zhang_TMLR_mm-cot and UnifiedQA Lu_NeurIPS22_unifiedqa are the pioneering works introducing the concept of CoT prompting into the multi-modal reasoning of LVLMs. Successor works focus on improving the quality of $r$ via auxiliary VQA models Zheng_NeurIPS23_ddcot and knowledge base retrieval Mondal_AAAI24_kam-cot . Although these works successfully elicited LVLM reasoning, their reliance on additional training and/or auxiliary resources (e.g., VQA models and knowledge bases) limits broader applicability. To address this limitation, CCoT Mitra_CVPR24_CCoT enhanced rationale quality via structured scene graphs in JSON format generated from LVLMs’ zero-shot reasoning without additional training. Most existing works focus on improving $r$ in Eq. (5) and assume that accurate $r$ ensures better next token prediction with $p_{\theta}(y_{i}|\bm{y}_{<i},x,r,q)$. In this regard, our work is orthogonal to them because we investigate the reliability of $p_{\theta}(y_{i}|\bm{y}_{<i},x,r,q)$ as discussed in the next section. This work focuses on training-free CoT reasoning for maintaining simplicity and generalization of pre-trained LVLMs, but our findings can extend to any CoT methods in a plug-and-play manner.
2.3 Motivating Experiments
We show our motivation through preliminary experiments asking a simple question: Do LVLMs perform CoT reasoning grounded on intermediate rationales? Specifically, we assess how much $r$ and its contents contribute to the output sequence decoded by $p_{\theta}(y_{i}|\bm{y}_{<i},x,r,q)$ through two preliminary experiments: (i) measuring the contribution scores to output token predictions for input token groups corresponding to the image $x$, rationale $r$, and query $q$, and (ii) replacing rationale $r$ with $r^{\prime}$ from another $(x^{\prime},q^{\prime})$ pair. Experiment (i) aims to evaluate how much $r$ contributes to output $\bm{y}$, and (ii) checks if LVLMs leverage the content of $r$ associated with $(x,q)$. We used Gemma-3-4B/12B Google_2025_gemma3 as the LVLM and GQA Hudson_CVPR19_gqa as the evaluation dataset. For the rationale generation, we queried LVLMs with Eq. (5) to generate text description rationales and scene graphs with the prompts of CCoT Mitra_CVPR24_CCoT ; hereinafter, we refer to the reasoning with text descriptions as CoT and that with scene graphs as CCoT. See Appendix A for details.
@@ 2.3 Motivating Experiments
Weakened influence of rationales. Here, we measure attention contribution scores Kobayashi_EMNLP20_attention_contribution ; Basu_NeurIPS24_understanding_info_lvlm ; Kang_ICLR25_visual_attention_sink to quantify input token $(x,r,q)$ contributions to output $\bm{y}$. Ideally, CoT prediction should derive substantial contributions from both image and rationale tokens. The contribution score from $i$-th token to $j$-th token at the $l$-th layer and $h$-th head is computed by $\|\alpha^{l,h}_{i,j}\bm{z}^{l-1}_{j}\bm{W}^{l,h}_{OV}\|$, where $\alpha^{l,h}_{i,j}$ is attention score, $\bm{z}^{l-1}_{j}$ is the output of the previous layer, and $\bm{W}^{l,h}_{OV}$ is the output projection matrix in transformer-based LVLMs. We describe more details of the evaluation protocol in Appendix A. Figure 1 shows the attention contributions in the middle layer of Gemma-3-12B for each case of input for decoding, where the scores are computed for each token corresponding to $x$, $r$, and $q$, respectively, and displayed as percentages of the overall contribution score for each input type. While image and rationales contribute largely when conditioned individually (i.e., $(x,q)$, $(r_{\mathrm{CoT}},q)$, and $(r_{\mathrm{CCoT}},q)$), in multi-modal CoT (i.e., $(x,r_{\mathrm{CoT}},q)$, and $(x,r_{\mathrm{CCoT}},q)$), image tokens dominate and rationale contribution is remarkably reduced.
@@
Lack of grounding on rationale contents. Next, we analyze rationale content effects by replacing original rationale $r$ with an irrelevant rationale $r^{\prime}$ generated from another random $(x^{\prime},q^{\prime})$ pair. We expect that intervening by $r^{\prime}$ will largely degrade the performance. Table 1 shows the test accuracy on GQA for the combinations of decoding strategy and input for decoding. While CoT and CCoT with $(x,r,q)$ improved the baseline performance with Gemma-3-4B, they degraded the performance with Gemma-3-12B. More importantly, in Table 1, CoT/CCoT with the irrelevant rationale $r^{\prime}$ via $(x,r^{\prime},q)$ performed similarity to using original $r$, indicating that LVLMs ignore the contents of $r$ and the performance gain comes from other causes. Conversely, decoding with $(r^{\prime},q)$ largely dropped the scores from that with $(r,q)$, implying that LVLMs use $r$ when $x$ is absent. This phenomenon could potentially be attributed to several factors, including position bias in the LLM backbones Zhang_NeurIPS24_pos_bias ; Zhang_EMNLP24_instruct_pos_bias , attention sink Xiao_ICLR24_attention_sink ; Gu_ICLR25_when_attention_sink_emerges ; Kang_ICLR25_visual_attention_sink , and/or overfitting to visual instruction tuning Li_EMNLP23_evaluating_lvlm ; He_ICLR25_analyzing_FGVR_lvlm . Since our focus is training-free CoT in off-the-shelf LVLMs, we leave a deeper analysis for future work.
@@
In summary, the two experiments demonstrate that LVLMs are less likely to leverage rationale $r$ mechanistically and semantically in the multi-modal CoT prompting. This challenge is important because CoT prompting is generally expected to enhance not only task performance but also faithfulness and interpretability Lanham_2023_measuring_faithfulness_cot ; Paul_EMNLP24_measuring_improving_faithfulness_cot ; Jie_NAACL24_interpretability_cot . This motivates our new decoding strategy to ground the inference on $r$.
@@ 3.1 Multi-modal CoT as a Reward Maximization on Rationale
As shown in Section 2.3, conventional multi-modal CoT prompting via $p_{\theta}(y_{i}|\bm{y}_{<i},x,r,q)$ often fails to effectively leverage rationale $r$ for output predictions. We also observed that the prediction with $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$ relies on $r$, but the performance is suboptimal. We aim for a next token distribution more effective than both $p(y_{i}|\bm{y}_{<i},x,r,q)$ and $p(y_{i}|\bm{y}_{<i},r,q)$. To this end, our core idea is decoupling the next token probability into distinct image-conditional $p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$ and rationale-conditional $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$ and separately enhancing rationale grounding. Enhancing rationale grounding indeed corresponds to maximizing the log-likelihood of $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$; we refer to this log-likelihood as rationale-grounding reward. Therefore, instead of $p_{\theta}(y_{i}|\bm{y}_{<i},x,r,q)$, we introduce a new next token distribution $\pi$ maximizing the rationale-grounding reward based on a KL-constrained reward maximization Schulman_ICML15_TRPO ; Schulman_ICML15_PPO as follows:

### v2

@@ Abstract
Figure 1: Rationale-Enhanced Decoding (RED). Existing multi-modal chain-of-thought (CoT) prompting by large vision language models (LVLMs) is a two-step generation of rationale and final output. It often focuses on input images and overlooks intermediate rationales in the final output generation. Our rationale-enhanced decoding (RED) addresses this issue by decoupling the image and rationale in decoding, and combining them at the logit level to provably ensure grounding outputs on the rationale.
@@ 1 Introduction
Recent large language model (LLM) advancements exhibit impressive complex reasoning capabilities [1, 54, 4]. This progress now extends beyond single-modal text-to-text reasoning. The integration of pre-trained vision encoders, like CLIP [43], with LLMs has led to the development of large vision-language models (LVLMs) [32, 10, 76, 8, 3], enabling more complex multi-modal reasoning.
Chain-of-thought (CoT) prompting is a key factor in LLM reasoning abilities [60, 23]. CoT-prompted models first generate intermediate reasoning steps, termed rationales, then incorporate them into the context to produce the final query response. CoT facilitates models to understand queries deeply and promotes logical and coherent responses [56, 45, 59]. This principle has also been applied to multi-modal reasoning in LVLMs [71, 34, 73, 16, 38, 37, 74, 72]. In CoT for LVLMs, models generate intermediate rationales from the input image and the text query. These rationales, with the original inputs, are then used for grounded image-text-to-text reasoning. Recent research also explores generating structured rationales (e.g., scene graphs) to enhance LVLMs’ capabilities like spatial reasoning [37, 72]. Thus, CoT in LVLMs is widely assumed to be beneficial, offering multi-modally grounded understanding for more accurate responses.
@@
A straightforward solution to make models force rationale grounding would be to fine-tune LVLMs on datasets containing image-query-rationale-answer tuples, as in [71, 73]. However, this approach demands expensive annotated datasets and additional training costs. Therefore, we focus on developing a plug-and-play decoding strategy for pre-trained LVLMs. We address the following primary research question: Can we enhance the LVLMs’ grounding capability on the rationales in CoT and improve performance by solely modifying the decoding strategy without additional training?
@@
In this paper, we propose rationale-enhanced decoding (RED), a novel decoding strategy designed to harmonize information from visual and rationale tokens in LVLMs without additional training. Our core idea is decoupling the next token probability into distinct image-conditional $p(y|x,q)$ and rationale-conditional $p(y|r,q)$ and separately enhancing rationale grounding ($x$: image, $r$: rationale, $q$: query, and $y$: output). Based on this idea, we re-formulate multi-modal CoT prompting as a KL-constrained reward maximization [46, 47]. This aims to maximize policy regarding the rationale-conditional log-likelihood $\log p(y|r,q)$ as the reward while staying close to the image-conditional likelihood $p(y|x,q)$. Solving this problem makes LVLMs explicitly ground on both image and rationale information in the next token prediction. Without additional training, RED provably yields the optimal solution of this maximization by composing the next-token distribution as $p(y|x,q)\times p(y|r,q)^{\lambda}$. Practically, RED is implemented as a simple weighted sum of the log-softmax logits of $p(y|x,q)$ and $p(y|r,q)$, allowing easy, training-free, and plug-and-play integration with existing LVLMs without architecture modification.
@@
We conduct extensive experiments comparing RED against standard CoT prompting and plug-and-play decoding methods for off-the-shelf LVLMs across multiple benchmark datasets and backbone LVLMs. Our results demonstrate that RED consistently and significantly improves reasoning performance. Furthermore, RED’s advantages are amplified with high-quality rationales (e.g., intervening with GPT-4[1]). Our findings validate the effectiveness and practicality of RED for enhancing CoT reasoning faithfulness and accuracy in LVLMs, opening avenues for more reliable, interpretable, rationale-grounded multi-modal systems.
2 Related Work
Large Vision Language Models (LVLMs).
Large vision language models (LVLMs) integrate visual encoders with LLMs, often via alignment training to represent images in the LLM input space [32, 10, 76, 8]. Despite remarkable multi-modal capabilities [25, 41, 39], LVLMs face challenges like poor visual recognition [69], object hallucination [31, 14, 62], and misalignment between image and text tokens [66, 57, 6]. Furthermore, as our preliminary experiments (Section 3.3) highlight, LVLMs also struggle to effectively leverage rationales in CoT reasoning. Previous solutions include preference/reward tuning [75, 52], improved visual instruction tuning [40, 7, 48, 9, 50], auxiliary model-enhanced decoding [64, 67, 51], plug-and-play decoding [36, 26, 58, 12], and CoT prompting [71, 73, 38, 16, 74, 37]. Our RED is categorized into a plug-and-play decoding for improving CoT prompting. While many CoT prompting methods focus on improving rationale generation via specific prompting or auxiliary model training (see Section 3.2), they often still rely on standard decoding with $p_{\theta}(y_{i}|\bm{y}_{<i},x,r,q)$ that may not leverage rationales. Our work diverges with a novel plug-and-play decoding strategy designed to enhance rationale utilization in CoT reasoning, without additional training or auxiliary models.
Plug-and-play Decoding Strategies for LVLMs.
Plug-and-play decoding strategies [36, 26, 58, 21, 12] are relevant as they operate at inference time without additional training like RED. These decoding strategies are mainly focused on mitigating object hallucination by contrastive decoding [29]. For instance, LCD [36] contrasts $p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$ with $p_{\theta}(y_{i}|\bm{y}_{<i},q)$ to mitigate the language prior effects; VCD [26] subtracts hallucinated predictions by contrasting $p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$ with $p_{\theta}(y_{i}|\bm{y}_{<i},x^{\prime},q)$ ($x^{\prime}$ is the corrupted image). Thus, these decoding strategies basically aim to modulate image-conditional probability $p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$ to mitigate object hallucination. However, these methods do not ensure LVLMs faithfully use CoT rationales; they refine image-conditional output, not harmonize it for CoT reasoning. Therefore, RED is complementary to these decoding strategies because it grounds predictions by multiplying distinct image-conditional $p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$, which is potentially pre-modulated by other methods, and rationale-conditional $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$. This allows combining RED with other plug-and-play methods for synergistic benefits of better hallucination mitigation and rationale grounding.
@@ 3 Preliminaries
3.1 Next-token Prediction in LVLMs
@@
$p_{\theta}(y_{i}|\bm{y}_{<i},x,q)=\operatorname{softmax}(\operatorname{logits}_{\theta}(y_{i}|\bm{y}_{<i},x,q))$ (2)
@@
Each token $y_{i}$ is generated from $p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$ via a decoding strategy such as greedy decoding, i.e., $y_{i}=\mathop{\rm arg~max}\limits_{w\in\mathcal{V}}p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$. We define $\operatorname{generate(\cdot)}$ as a utility function for an arbitrary decoding strategy:
@@ 3.2 Multi-modal Chain-of-Thought Prompting
Inspired by single modal CoT prompting [60], multi-modal CoT aims to enhance LVLM reasoning by incorporating the intermediate rationale generation during decoding [71, 73, 38, 37]. In general, multi-modal CoT involves two reasoning steps: (i) rationale generation and (ii) output generation. First, LVLMs generate a rationale $r$ from input image $x$, query prompt $q$ with instruction prompt, e.g., “Given the image, generate the rationale for answering the question”, as follows.
@@
In this line of work, MM-CoT [71] and UnifiedQA [34] are the pioneering works introducing the concept of CoT prompting into the multi-modal reasoning of LVLMs. Successor works focus on improving the quality of $r$ via auxiliary VQA models [73] and knowledge base retrieval [38]. Although these works successfully elicited LVLM reasoning, their reliance on additional training and/or auxiliary resources (e.g., VQA models and knowledge bases) limits broader applicability. To address this limitation, CCoT [37] enhanced rationale quality via structured scene graphs in JSON format generated from LVLMs’ zero-shot reasoning without additional training. Most existing works focus on improving $r$ in Eq. (5) and assume that accurate $r$ ensures better next token prediction with $p_{\theta}(y_{i}|\bm{y}_{<i},x,r,q)$. In this regard, our work is orthogonal to them because we investigate the reliability of $p_{\theta}(y_{i}|\bm{y}_{<i},x,r,q)$ as discussed in the next section. This work focuses on training-free CoT reasoning for maintaining simplicity and generalization of pre-trained LVLMs, but our findings can extend to any CoT methods in a plug-and-play manner.
@@
3.3 Motivating Experiments
We show our motivation through preliminary experiments asking a simple question: Do LVLMs perform CoT reasoning grounded on intermediate rationales? Specifically, we assess how much $r$ and its contents contribute to the output sequence decoded by $p_{\theta}(y_{i}|\bm{y}_{<i},x,r,q)$ through two preliminary experiments: (i) measuring the contribution scores to output token predictions for input token groups corresponding to the image $x$, rationale $r$, and query $q$, and (ii) replacing rationale $r$ with $r^{\prime}$ from another $(x^{\prime},q^{\prime})$ pair. Experiment (i) aims to evaluate how much $r$ contributes to output $\bm{y}$, and (ii) checks if LVLMs leverage the content of $r$ associated with $(x,q)$. We used Gemma-3-4B/12B [55] as the LVLM and GQA [18] as the evaluation dataset. For the rationale generation, we queried LVLMs with Eq. (5) to generate text description rationales and scene graphs with the prompts of CCoT [37]; hereinafter, we refer to the reasoning with text descriptions as CoT and that with scene graphs as CCoT. See Appendix for details.
@@ 3.3 Motivating Experiments
Weakened influence of rationales. Here, we measure attention contribution scores [22, 5, 20] to quantify input token $(x,r,q)$ contributions to output $\bm{y}$. Ideally, CoT prediction should derive substantial contributions from both image and rationale tokens. The contribution score from $i$-th token to $j$-th token at the $l$-th layer and $h$-th head is computed by $\|\alpha^{l,h}_{i,j}\bm{z}^{l-1}_{j}\bm{W}^{l,h}_{OV}\|$, where $\alpha^{l,h}_{i,j}$ is attention score, $\bm{z}^{l-1}_{j}$ is the output of the previous layer, and $\bm{W}^{l,h}_{OV}$ is the output projection matrix in transformer-based LVLMs. We describe more details of the evaluation protocol in Appendix. Figure 2 shows the attention contributions in the middle layer of Gemma-3-12B for each case of input for decoding, where the scores are computed for each token corresponding to $x$, $r$, and $q$, respectively, and displayed as percentages of the overall contribution score for each input type  [^1] We analyzed a middle layer because prior work shows it is where LVLMs primarily perform multi-modal fusion to integrate vision and language [39]. This allows us to observe the model’s attention balance between the image and rationale during this crucial integration process, before the representation becomes overly specialized for the final output task.. While image and rationales contribute largely when conditioned individually (i.e., $(x,q)$, $(r_{\mathrm{CoT}},q)$, and $(r_{\mathrm{CCoT}},q)$), in multi-modal CoT (i.e., $(x,r_{\mathrm{CoT}},q)$, and $(x,r_{\mathrm{CCoT}},q)$), image tokens dominate and rationale contribution is remarkably reduced.
@@
Lack of grounding on rationale contents. Next, we analyze rationale content effects by replacing original rationale $r$ with an irrelevant rationale $r^{\prime}$ generated from another random $(x^{\prime},q^{\prime})$ pair. We expect that intervening by $r^{\prime}$ will largely degrade the performance. Table 1 shows the test accuracy on GQA for the combinations of decoding strategy and input for decoding. While CoT and CCoT with $(x,r,q)$ improved the baseline performance with Gemma-3-4B, they degraded the performance with Gemma-3-12B. More importantly, in Table 1, CoT/CCoT with the irrelevant rationale $r^{\prime}$ via $(x,r^{\prime},q)$ performed similarity to using original $r$, indicating that LVLMs ignore the contents of $r$ and the performance gain comes from other causes. Conversely, decoding with $(r^{\prime},q)$ largely dropped the scores from that with $(r,q)$, implying that LVLMs use $r$ when $x$ is absent. This phenomenon could potentially be attributed to several factors, including position bias in the LLM backbones [70, 68], attention sink [63, 13, 20], and/or overfitting to visual instruction tuning [31, 15]. Since our focus is training-free CoT in off-the-shelf LVLMs, we leave a deeper analysis for future work.
@@
In summary, the two experiments demonstrate that LVLMs are less likely to leverage rationale $r$ mechanistically and semantically in the multi-modal CoT prompting. This challenge is important because CoT prompting is generally expected to enhance not only task performance but also faithfulness and interpretability [24, 42, 19]. This motivates our new decoding strategy to ground the inference on $r$.
@@ 4.1 Multi-modal CoT as a Reward Maximization on Rationale
As shown in Section 3.3, conventional multi-modal CoT prompting via $p_{\theta}(y_{i}|\bm{y}_{<i},x,r,q)$ often fails to effectively leverage rationale $r$ for output predictions. We also observed that the prediction with $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$ relies on $r$, but the performance is suboptimal. We aim for a next token distribution more effective than both $p(y_{i}|\bm{y}_{<i},x,r,q)$ and $p(y_{i}|\bm{y}_{<i},r,q)$. To this end, our core idea is decoupling the next token probability into distinct image-conditional $p_{\theta}(y_{i}|\bm{y}_{<i},x,q)$ and rationale-conditional $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$ and separately enhancing rationale grounding. Enhancing rationale grounding indeed corresponds to maximizing the log-likelihood of $p_{\theta}(y_{i}|\bm{y}_{<i},r,q)$; we refer to this log-likelihood as rationale-grounding reward. Therefore, instead of $p_{\theta}(y_{i}|\bm{y}_{<i},x,r,q)$, we introduce a new next token distribution $\pi$ maximizing the rationale-grounding reward based on a KL-constrained reward maximization [46, 47] as follows: